In the following reaction the cyanoacetate is firs
Answer #1
Concepts and reason

Condensation reaction: The reaction in which two small compounds react together and form larger molecules. During condensation reaction, elimination of small molecules such as water, methanol or HClHCl will take place.

Knoevenagel condensation: It is nucleophilic addition of an α{\rm{\alpha }} -hydrogen compound to the carbonyl carbon, followed by deprotonation. The obtained product is α,β{\rm{\alpha ,\beta }} - unsaturated ketone. The α{\rm{\alpha }} -hydrogen is the hydrogen of the carbon adjacent to the functional group.

αHydrogen{\bf{\alpha }}\,{\bf{ - }}\,{\bf{Hydrogen}} : The hydrogen bonded to the carbon(s) that is/are adjacent to a functional group, is termed as the αcarbon{\rm{\alpha - carbon}} , hydrogen atom bonded to such carbon atom is called as αhydrogen{\rm{\alpha }}\,{\rm{ - }}\,{\rm{hydrogen}} . In case of electron withdrawing group αhydrogen{\rm{\alpha }}\,{\rm{ - }}\,{\rm{hydrogen}} are acidic in nature, due to the partial positive charge created in αcarbon{\rm{\alpha - carbon}} atom.

Fundamentals

Representative example Knoevenagel condensation

base
R-CHO
+
CH2(COOR)2
R-CHEC(COOR)2
-H2O
CO2 |
он
R-CHECH(COOH)

When cynoacetate is first treated with methoxide, then slowly it is added to the aldehyde. The product will be formed as follows:

0
—
—
y
СН,ОН

The mechanism involved for the formation of product is given below:تهران |
A-
CH, OH
CH,0
CH,0
CH,
OH

Elimination of water molecule:

НО
о“Н-
H—
HC

Ans:

When cynoacetate reacts with methoxy ion, then slowly it is added to aldehyde and the product that will be formed as follows:

CH3OH